a gun shoots bullets that leave the muzzle at 250 m/s. If the bullet is to hit a target 100m away at the level of the muzzle, the gun must be aimed. How far above the target is this point?
Answer #1

Solution) u = 250 m/s

X = 100 m

H = ?

We have

X = ut

t = (X/u)

t = (100/250)

t = 0.4 s

H = Vt + (1/2)gt^2

V = 0 ( along vertical direction velocity is zero)

H = (1/2)(g)(t^2)

H = (1/2)(9.8)(0.4^2)

H = 0.784 m